OpenStudy (anonymous):

will we use chain rule when ever we have to find derivative of a function with respect to itself ?

OpenStudy (abb0t):

for example, if you have: \(\large sin^{200}(8x)\) this is the same thing as saying \(\large [sin(8x)]^{200}\) in which case you are using chain rule. The derivative of the outside FIRST mutiplied by derivative of the inside. So your outside is the 200, your inside is ur sin, and then inside of ur sine, you have (8x) So you would get: \( large (200 \times [sin(8x)]^{200-1} \times cos(8x) \times 8\) you see how i did that?

OpenStudy (anonymous):

yeah you used d\dx(f^n)= nf^n-1 * d\dx(f)

OpenStudy (abb0t):

yes

OpenStudy (anonymous):

but what will be derivative of tanx w.r.t to tan x ? it will be 1 by using chain rule which is my question ?

OpenStudy (abb0t):

so \(tan(tan(x))\)??? sme as chain rule. derivative of outside tan times derivative of inside tangent.

OpenStudy (abb0t):

its the same with \(cos(cos(cos(cos(cos(cos(x)))))))\)